# The Intelligible and Effective Graph Neural Additive Networks

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as the predominant approach for learning over graph-structured data. However, most GNNs operate as black-box models and require post-hoc explanations, which may not suffice in high-stakes scenarios where transparency is crucial. In this paper, we present a GNN that is interpretable by design. Our model, Graph Neural Additive Network (GNAN), is a novel extension of the interpretable class of Generalized Additive Models, and can be visualized and fully understood by humans. GNAN is designed to be fully interpretable, allowing both global and local explanations at the feature and graph levels through direct visualization of the model. These visualizations describe the exact way the model uses the relationships between the target variable, the features, and the graph. We demonstrate the intelligibility of GNANs in a series of examples on different tasks and datasets. In addition, we show that the accuracy of GNAN is on par with black-box GNNs, making it suitable for critical applications where transparency is essential, alongside high accuracy.
@@ 1 Introduction
In many domains, ranging from biology to fraud detection, Artificial Intelligence (AI) is applied to data with graph structure. Neural Networks, and specifically Graph Neural Networks (GNNs), have emerged as the predominant approach in these applications (see, for example, Zhou et al. [1]). While GNNs demonstrate high accuracy, in terms of the correctness of their predictions, they often function as black-box models; thus, their decision-making processes are opaque. Transparency is vital for assessing potential biases or safety risks and is particularly critical in high-stakes areas such as criminal justice, healthcare, and finance, where decisions significantly impact individuals’ lives. In such contexts, interpretable models, despite sometimes being less accurate, may be preferred over complex black-box models [2]. Furthermore, the transparency of automated decision making processes is increasingly becoming a legal mandate. While there is ongoing debate over whether the European Union’s General Data Protection Regulation (GDPR) implies a “right to explanation” [3, 4], the proposed European AI Act explicitly addresses this issue, stating that “To address concerns related to opacity and complexity of certain AI systems and help deployers to fulfill their obligations under this regulation, transparency should be required for high-risk AI systems before they are placed on the market or put into service” [5].
@@
Methods for model explainability or interpretability can be categorized into local and global types. Local methods, such as SHAP [12] and LIME [13], elucidate individual predictions made by a model, whereas global methods, such as feature-importance [14] and partial dependence plots [15], provide holistic insights about the model, i.e., explain the overarching logic of the model decision making [16]. However, it has been noted that local explainability methods may not consistently align with their global counterparts [17]. Moreover, local explanations may be inadequate for verifying fairness and other risks [8].
@@ 2 Related work
Generalized Additive Models (GAMs) are a class of statistical models that build upon generalized linear models by incorporating non-linear functions for each variable while maintaining additivity [19, 20, 21]. Essentially, GAMs model the expected value of the target variable as a sum of univariate functions of the features. Formally, in GAMs, a prediction for an input $x$ is computed by $\sigma\left(\sum f_{k}\left(x_{k}\right)\right)$ where $\sigma$ is a predefined activation function, such as the sigmoid [^1] In the context of Generalized Linear Models (GLMs) $\sigma$ can be though of as the inverse of the link-function., and the $f_{k}$’s are shape functions learned during the training process. This approach extends generalized linear models, in which predictions are computed by $\sigma\left(\sum w_{k}x_{k}\right)$ where $w$ is a learned weight vector.
@@ 3 Graph Neural Additive Networks
$\sigma\left(\sum_{k=1}^{d}[\mathbf{h}]_{k}\right)~{}~{}~{}.$
@@
From the definitions provided above, it follows that the entire model can be represented with just a few figures, thus providing global interpretability. For local explanations, it is feasible to examine the contribution of each feature and each node to the predictions. Furthermore, consider the following graph representation:
$[\mathbf{h}]_{k}=\sum_{i=1}^{N}[\mathbf{h}_{i}]_{k}=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\cdot f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)$
$=\sum_{j=1}^{N}f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)\sum_{i=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)~{}~{}~{}.$
This mathematical formulation reveals how each node and feature contribute to the overall graph representation. The elements in the first equation convey the effect of nodes, while the elements of the second equation convey the influence of individual features across all nodes. Therefore, the model facilitates a detailed understanding of local behavior from multiple perspectives.
@@ 4 Inteligibility
To visualize the contribution of a feature value at a specific distance, we employ a heatmap for each class, evaluating the products between the outputs of the feature function over the input range ($[0,1]$) and the output of the corresponding distance function. Figure 5 exemplifies this visualization technique with the ’children’ feature. It is insightful to observe that the presence of the word ’children’ influences the predictions differently across the diabetes types. The model has learned that papers concerning type 1 diabetes seldom mention ’children’, nor do related papers. In contrast, the term frequently appears in the context of gestational diabetes. Further examples of GNAN visualizations are presented in the Appendix.
@@ 5 Empirical evaluation
For all tasks, we used existing splits, protocols, and metrics, as commonly used in the literature for each dataset. The complete protocols for each dataset are given in detail in the Appendix. The metrics we report are: For Cornell, Cora, Citeseer, PubMed, ogb-arxiv, Mutagenicity, PTC, NCI, and Proteins, we report accuracy. For $\mu$, $\alpha$, $\alpha_{HOMO}$ we report MAE. For Tolokers we report ROC-AUC. For the node labeling tasks, we used the pre-defined splits in the data and followed the common protocols for each dataset. The results are an average of the test set using $5$ or $10$ random seeds. For the Proteins and NCI1 tasks, we followed the splits and the nested-cross-validation protocol from [58]. The final reported result on these datasets is an average of 30 runs (10-folds and 3 random seeds). For NCI1 and PTC we followed the splits and protocol from [39] and report the average accuracy and std of a 10-fold nested cross-validation.
@@ 6 Conclusion
In this work, we introduced the Graph Neural Additive Network (GNAN), a novel extension of the interpretable class of Generalized Additive Models, to accommodate graph data. GNAN is inherently interpretable, and provides both global and local explanations directly from its architecture, eliminating the need for post-hoc interpretations and thus enhancing transparency. Furthermore, GNAN demonstrates competitive performance with popular GNNs, showing that intelligibility does not necessarily entail a significant degradation in accuracy.
It is possible to enhance GNAN in several ways. To generate smooth shape functions, one could integrate techniques from the recently proposed Kolmogorov–Arnold Networks [46]. Increasing the capacity of GNAN is feasible by learning individual distance functions for each feature. Exploring reduced capacity is also intriguing, particularly in scenarios with many features, where it may be beneficial to employ regularization to limit the number of shape functions used. Additionally, applying these techniques to biological network datasets, such as protein interactions, could be a valuable tool to support scientific discoveries. These and other directions are left for future studies.
@@ Appendix A Efficient GNAN implementation with tensor products
We formulate GNAM case of classification with $C$ classes. For regression, the exact formulation holds with $C=1$. For the sake of notation, we assume that every tensor that has its last dimension, $C$, is permuted to have the last dimension as its first dimension, without stating it explicitly. This is necessary to achieve a valid tensor multiplication.
GNAM linearly combines the outputs of at least $d+1$ neural networks $m$ and $\{f_{l}\}_{l=1}^{d}$.
$f_{l}:\mathbb{R}\rightarrow\mathbb{R}^{1\times C}$ is a feature-transformation function, and it acts on the $l^{\prime}th$ feature of some given node $i$, i.e., $f_{l}(x^{i}_{l})$
$m:\mathbb{R}\rightarrow\mathbb{R}^{1\times C}$ is a distance-transformation function, and it acts on distances between nodes, i.e., $m(dist(j,i))$.
Given $A$ , we build the distance matrix $D$ , where $D_{i,j}=dist(j,i)$
We denote with $M$ the matrix of the transformed distances that is outputted by applying $m$ to each entry of $D$, i.e., $M_{i,j}=m(dist(j,i))$, $M\in\mathbb{R}^{C\times N\times N}$.
We denote with $F$ the matrix of the transformed feature is outputted by applying the corresponding $f_{l}$ for feature $l$ of each node in the graph, i.e., $F_{il}=f_{l}(x_{l}^{i})$, $F\in\mathbb{R}^{C\times N\times d}$
Any GNAM, both for node and graph tasks, first computes the matrix $M\cdot F\in\mathbb{R}^{C\times N\times d}$
Then, depending on the task.
@@ A.1 Node Tasks
$\phi(i)=[M\otimes F\otimes\mathbbm{1}_{C\times d\times 1}]_{i}=\sum_{l=1}^{d}\sum_{j=1}^{N}m(dist(j,i))\otimes f_{l}(x_{l}^{j})$
@@ A.2 Graph Tasks
For graph classification, we set $m$ and $f$ to output a scalar and aggregate the transformed features over the nodes, i.e., the row of $M\cdot F$, to form a fixed-size vector of size $d$.
@@
Then, we can apply another readout NAM:
@@
Such that $\bar{F}$ is the transformed features using the function $\{\bar{f}_{l}\}_{l=1}^{d}$ such that $\bar{f}_{l}:\mathbb{R}\rightarrow\mathbb{R}^{1\times C}$
@@ Appendix B Extensions and ablations
Appendix C Additional explanations examples
@@
C.1 Additional PubMed heatmaps
In the main text, we presented the heatmaps for the ’children’ feature. Here we provide additional heatmaps for additional features: the ’fat’ feature and the ’young’ feature, as presented in Figures 6 and 7.
@@ C.1 Additional PubMed heatmaps
C.2 Tolokers - Binary classification with binary and continuous features
Tolokers dataset is based on data from the Toloka crowdsourcing platform. The nodes represent tolokers (workers) who have participated in at least one of 13 selected projects. An edge connects two tolokers if they have worked on the same task. The goal is to predict which tolokers have been banned in one of the projects. Node features are based on the worker’s profile information and task performance statistics. Each node in the graph is associated with 9 features. There are 4 continuous features in the range $[0,1]$ and 5 are binary features. Figure 8 shows the shape functions of the features learned by GNAN. Figure 9 shows the distance shape function learned by GNAN. In Figures 11 and 10 we present the heatmaps of the cross product of the shape functions and the distance function.
@@ D.2 Protocols
Following [64, 30, 29], for the Core, Citeseer and Pubmed datasets we tuned the parameters on the Cora dataset using the pre-defined splits from [64]. For all these datasets we report the test accuracies averaged over $5$ runs, using the parameters obtained from the best accuracy on the validation set of Cora.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have emerged as the predominant approach for learning over graph-structured data. However, most GNNs operate as black-box models and require post-hoc explanations, which may not suffice in high-stakes scenarios where transparency is crucial. In this paper, we present a GNN that is interpretable by design. Our model, Graph Neural Additive Network (GNAN), is a novel extension of the interpretable class of Generalized Additive Models, and can be visualized and fully understood by humans. GNAN is designed to be fully interpretable, offering both global and local explanations at the feature and graph levels through direct visualization of the model. These visualizations describe exactly how the model uses the relationships between the target variable, the features, and the graph. We demonstrate the intelligibility of GNANs in a series of examples on different tasks and datasets. In addition, we show that the accuracy of GNAN is on par with black-box GNNs, making it suitable for critical applications where transparency is essential, alongside high accuracy.
@@ 1 Introduction
In many domains, ranging from biology to fraud detection, Artificial Intelligence (AI) is applied to data with graph structure. Neural Networks, and specifically Graph Neural Networks (GNNs), have emerged as the predominant approach in these applications (see, for example, Zhou et al. [1]). While GNNs demonstrate high accuracy, in terms of the correctness of their predictions, they often function as black-box models; thus, their decision-making processes are opaque. Transparency is vital for assessing potential biases or safety risks, and is particularly critical in high-stakes areas such as criminal justice, healthcare, and finance, where decisions significantly impact individuals’ lives. In such contexts, interpretable models, despite sometimes being less accurate, may be preferred over complex black-box models [2]. Furthermore, the transparency of automated decision making processes is increasingly becoming a legal mandate. While there is ongoing debate over whether the European Union’s General Data Protection Regulation (GDPR) implies a “right to explanation” [3, 4], the proposed European AI Act explicitly addresses this issue, stating that “To address concerns related to opacity and complexity of certain AI systems and help deployers to fulfill their obligations under this regulation, transparency should be required for high-risk AI systems before they are placed on the market or put into service” [5].
@@
Methods for model explainability or interpretability can be categorized into local and global types. Local methods, such as SHAP [12] and LIME [13], elucidate individual predictions made by a model, whereas global methods, such as feature-importance [14] and partial dependence plots [15], provide holistic insights about the model, i.e., explain the overarching logic of the model decision-making [16]. However, it has been noted that local explainability methods may not consistently align with their global counterparts [17]. Moreover, local explanations may be inadequate for verifying fairness and other risks [8].
@@ 2 Related work
Generalized Additive Models (GAMs) are a class of statistical models that build upon generalized linear models by incorporating non-linear functions for each variable while maintaining additivity [19, 20, 21]. Essentially, GAMs model the expected value of the target variable as a sum of univariate functions of the features. Formally, in GAMs, a prediction for an input $\mathbf{x}$ is computed by $\sigma\left(\sum f_{k}\left(\mathbf{x}_{k}\right)\right)$ where $\sigma$ is a predefined activation function, such as the sigmoid [^1] In the context of Generalized Linear Models (GLMs) $\sigma$ can be though of as the inverse of the link-function., and the $f_{k}$’s are shape functions learned during the training process. This approach extends generalized linear models, in which predictions are computed by $\sigma\left(\sum\mathbf{w}_{k}\mathbf{x}_{k}\right)$ where $w$ is a learned weight vector.
@@ 3 Graph Neural Additive Networks
$\sigma\left(\sum_{k=1}^{d}[\mathbf{h}]_{k}\right)~{}~{}~{}.$ (1)
@@
From the definitions provided above, it follows that the entire model can be represented with just a few figures, thus providing global interpretability. For local explanations, it is feasible to examine the contribution of each feature and each node to the predictions. The following representation convey the influence of each node to the $k$’ts feature:
$[\mathbf{h}]_{k}=\sum_{i=1}^{N}[\mathbf{h}_{i}]_{k}=\sum_{j=1}^{N}f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)\sum_{i=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)~{}~{}~{}.$
Here $[\mathbf{h}_{i}]_{k}$ contains the influence of the $k$’th feature in node $i$ on the prediction. However, from the definition of $[\mathbf{h}_{i}]_{k}$ we see that it serves as a mediator for influences of all other nodes. Therefore, the influence of node $j$ on feature $k$ in the final graph representation can be obtained by:
$f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)\sum_{i=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)~{}~{}~{}.$ (2)
We can also extract the total contribution of each node $i$ to the prediction, by summing the contribution of the nodes over the features, as done in the input to $\sigma$ in Equation 1:
$\sum_{k=1}^{d}[\mathbf{h}]_{k}=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\sum_{k=1}^{d}f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)~{}~{}~{}.$
Therefore, the total contribution of node $i$ to the prediction is
$\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\sum_{k=1}^{d}f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)~{}~{}~{}.$ (3)
Overall, the model facilitates a detailed understanding of local behavior from multiple perspectives.
@@ 4 Inteligibility
To visualize the contribution of a feature value at a specific distance, we employ a heatmap for each class, evaluating the products between the outputs of the feature function over the input range ($[0,1]$) and the output of the corresponding distance function. Figure 5 exemplifies this visualization technique with the ’children’ feature. It is insightful to observe that the presence of the word ’children’ influences the predictions differently across the diabetes types. The model has learned that papers concerning type 1 diabetes seldom mention ’children’, nor do related papers. In contrast, the term frequently appears in the context of gestational diabetes.
It is possible to construct confidence intervals for GAMs using the bootstrap method[49, 50]. We present such example with additional visualization examples in the appendix.
4.1 Local Explanations
Up to this point, we have demonstrated how to visualize GNAN for providing global explanations. These explanations offer a comprehensive visualization of the entire model. Now, we shift our focus to local explanations, i.e., those relevant to particular examples of interest. To illustrate this, we employ the same Mutagenicity model that was visualized globally to explain specific samples within the data. Using Equation  3 , we compute the importance of each node and visualize two molecules from the dataset, where the area of each node corresponds to its importance.
Figure LABEL:fig:local presents two such examples. In Figure LABEL:fig:carbon_ring the carbon (red) atoms play the most significant role, and a carbon ring (red cycle) is highlighted. In Figure LABEL:fig:no2, a group of $NO_{2}$ (grey and green subgraph) is shown to be relatively important for predicting the molecule as mutagenic. Both carbon rings and $NO_{2}$ groups are well-known for their mutagenic effects [51], making them frequently discussed in the literature on explainable GNNs [39, 41].
@@ 5 Empirical evaluation
For all tasks, we used existing splits, protocols, and metrics, as commonly used in the literature for each dataset. The complete protocols for each dataset are given in detail in the Appendix. The metrics we report are: For Cornell, Cora, Citeseer, PubMed, ogb-arxiv, Mutagenicity, PTC, NCI, and Proteins, we report accuracy. For $\mu$, $\alpha$ and $\alpha_{HOMO}$ we report MAE. For Tolokers we report ROC-AUC. For the node labeling tasks, we used the pre-defined splits in the data and followed the common protocols for each dataset. The results are an average of the test set using $5$ or $10$ random seeds. For the Proteins and NCI1 tasks, we followed the splits and the nested-cross-validation protocol from [61]. The final reported result on these datasets is an average of 30 runs (10-folds and 3 random seeds). For NCI1 and PTC we followed the splits and protocol from [39] and report the average accuracy and std of a 10-fold nested cross-validation.
@@ 6 Conclusion
In this work, we introduced the Graph Neural Additive Network (GNAN), a novel extension of the interpretable class of Generalized Additive Models, to accommodate graph data. GNAN is inherently interpretable, and provides both global and local explanations directly from its architecture, eliminating the need for post-hoc interpretations. This direct interpretability enhances the transparency of the model and is particularly useful in high-stakes applications where understanding model decisions is crucial. Furthermore, GNAN demonstrates competitive performance with popular GNNs, showing that intelligibility does not necessarily entail a significant degradation in accuracy.
It is possible to enhance GNAN in several ways. To generate smooth shape functions, one could integrate techniques from the recently proposed Kolmogorov–Arnold Networks [46] or adaptive activations for graphs [64]. Increasing the capacity of GNAN is feasible by learning individual distance functions for each feature. Exploring reduced capacity is also intriguing, particularly in scenarios with many features, where it may be beneficial to employ regularization to limit the number of shape functions used. Additionally, applying these techniques to biological network datasets, such as protein interactions, could be a valuable tool to support scientific discoveries. These and other directions are left for future studies.
@@ Appendix A Efficient GNAN implementation with tensor products
We formulate GNAN case of classification with $C$ classes. For regression, the exact formulation holds with $C=1$. We use the same notation as in the main text, only now the output of the feature and distance function is of dimension $1\times C$,
For the sake of notation, we assume that every tensor that its last dimension is of size $C$, is permuted to have the last dimension as its first dimension, without stating it explicitly. This is necessary to achieve a valid tensor multiplication.
We denote with $M$ the matrix of the transformed distances that is outputted by applying $\rho$, i.e., $M_{i,j}=\rho(\frac{1}{1+dist(j,i)})$, $M\in\mathbb{R}^{C\times N\times N}$.
We denote with $F$ the matrix of the transformed feature is outputted by applying the corresponding $f_{k}$ for feature $l$ of each node in the graph, i.e., $F_{ik}=f_{k}(x_{l}^{i})$, $F\in\mathbb{R}^{C\times N\times d}$
Both for node and graph tasks, we first computes the matrix $M\cdot F\in\mathbb{R}^{C\times N\times d}$
The rest of the computation then depends on the task.
@@ A.1 Node Tasks
$\phi(i)=[M\otimes F\otimes\mathbbm{1}_{C\times d\times 1}]_{i}=\sum_{k=1}^{d}\sum_{j=1}^{N}\rho(\frac{1}{1+dist(j,i)})\otimes f_{k}(\left[\mathbf{x}_{j}\right]_{k})$
@@ A.2 Graph Tasks
For graph classification, we set $\rho$ and $f$ to output a scalar and aggregate the transformed features over the nodes, i.e., the row of $M\cdot F$, to form a fixed-size vector of size $d$.
@@
Then, we can apply another readout NAM [26]:
@@
Such that $\bar{F}$ is the transformed features using the function $\{\bar{f}_{k}\}_{k=1}^{d}$ such that $\bar{f}_{k}:\mathbb{R}\rightarrow\mathbb{R}^{1\times C}$
@@ Appendix B Extensions and ablations
Appendix C Additional explanation examples
@@
C.1 Confidence Intervals
In our main paper, we discussed the construction of confidence intervals for Generalized Additive Models (GAMs) using the bootstrap method. Here, we provide a specific example using the Mutagenicity dataset.
We computed 95% confidence intervals by applying the bootstrap method with 200 resamples. This involved resampling the original dataset with replacement 200 times and calculating the statistic of interest for each resample. The resulting bootstrap estimates were sorted, and the 2.5th and 97.5th percentiles were taken as the lower and upper bounds of the confidence interval, respectively. Figure 7 presents the distance function with its confidence intervals.
Figure 7:
C.2 Additional PubMed heatmaps
In the main text, we presented the heatmaps for the ’children’ feature. Here we provide additional heatmaps for additional features: the ’fat’ feature and the ’young’ feature, as presented in Figures 8 and 9.
@@ C.2 Additional PubMed heatmaps
C.3 Tolokers - Binary classification with binary and continuous features
The Tolokers dataset is based on data from the Toloka crowdsourcing platform. The nodes represent tolokers (workers) who have participated in at least one of 13 selected projects. An edge connects two tolokers if they have worked on the same task. The goal is to predict which tolokers have been banned in one of the projects. Node features are based on the worker’s profile information and task performance statistics. Each node in the graph is associated with nine features. There are 4 continuous features in the range $[0,1]$ and 5 are binary features. Figure 10 shows the shape functions of the features learned by GNAN. Figure 11 shows the distance shape function learned by GNAN. In Figures 13 and 12 we present the heatmaps of the cross product of the shape functions and the distance function.
@@ D.2 Protocols
Following [68, 30, 29], for the Core, Citeseer and Pubmed datasets we tuned the parameters on the Cora dataset using the pre-defined splits from [68]. For all these datasets we report the test accuracies averaged over $5$ runs, using the parameters obtained from the best accuracy on the validation set of Cora.For Cora and Citeseer, we followed the schemes used in [69] where the used graph is the largest connected component, to increase the stability of the evaluated models.